Here are some balance of payments data (without pluses and minuses): Exports of goods, 100 Imports of goods, 125 Service exports, 90 Service imports, 80 Income receipts from abroad, 110 Income payments to foreigners, 150 Increase in home country’s ownership of assets abroad, 160 Increase in foreign ownership of assets in home country, 200 Increase in home reserve assets, 30 Increase in foreign reserve assets, 35 Assuming that unilateral transfers equal zero, find net exports, the current account balance, the capital and financial account balance, the official settlements balance, and the statistical discrepancy. (Note: The increase in home reserve assets of 30 is included in the increase in the home country’s ownership of assets abroad of 160, and the increase in foreign reserve assets of 35 is included in the increase in foreign ownership of assets in the home country of 200.)


1
Expert's answer
2021-10-26T09:33:35-0400

Solution:

Net exports - Goods = Exports – Imports = 100 – 125 = -25

Net exports – Services = Exports – Imports = 90 – 80 = 10

Total net exports = -25 + 10 = -15

Net income from abroad = 110 – 150 = -40

Current account balance = -25 + 10 – 40 = -55

Capital and financial account balance = -160 + 200 + 15 = 55

Net holdings of reserve assets = -100 + 125 – 90 + 80 – 110 + 150 + 160 – 200 = 15

 

Statistical discrepancy (SD): CA + KFA + SD = 0

–55 + 40 + SD = 0

Statistical discrepancy (SD) = 15